OpenStudy (anonymous):

Please help me on this problem. Thanks I have 4 cases of bottled water. There are 24 bottles in each case. How many bottles do you have? I need to find the unit rate, it tell me how many cases. I did the math and every answer i give is wrong

OpenStudy (tyteen4a03):

Well the answer is 4 * 24. Not sure what "unit rate" you're talking about.

OpenStudy (anonymous):

i typed 96 but it said it's wrong.

OpenStudy (anonymous):

the unit rate is 4/96 = 2/48 = 1/24 (I think)

OpenStudy (anonymous):

i understand.....thanks for explaining that to me.